OpenStudy (anonymous):

find the geometric mean of 45,_,_,20 with solution please

OpenStudy (sirm3d):

let r be the common ratio

OpenStudy (sirm3d):

the next term is obtained by mutiplying the previous term by r, in this case, what follows 45 is 45r.

OpenStudy (sirm3d):

the previous term is equal to the next term divided r. therefore the term before 20 is 20/r

OpenStudy (sirm3d):

the geometric mean is \[\large \sqrt[4]{45\cdot 45r \cdot \frac{20}{r}\cdot 20}=\sqrt[4]{45^2\cdot 20^2}=30\]